# Quantum-centric Supercomputing for Materials Science: A Perspective on Challenges and Future Directions

### v1

@@ II.2.2 Linear Combination of Unitaries
Another approach to the simulation of $e^{-iHt}$ involves block-encoding of the Hamiltonian as a part of a larger unitary, $U(H)$. One of the most typical block-encodings is accomplished by writing $H$ as a linear combination of unitaries, $H=\sum_{n}c_{n}U_{n}$, and building $U(H)$ using $U_{P}=\sum_{n}\mathcal{P}_{n}\otimes U_{n}$, where $\mathcal{P}_{n}$’s form a complete set of orthogonal projectors, $\mathcal{P}_{n}\mathcal{P}_{n^{\prime}}=\mathcal{P}_{n}\delta_{nn^{\prime}}$ built in the ancilla qubit space. The computational cost of the LCU-based algorithms depends on the $l1$-norm of $c_{n}$ coefficient vector, $\lambda=\sum_{n}|c_{n}|$ [159]. Thus, much effort was directed to reducing the $l1$-norm by various selections of $U_{n}$ operators [160, 161, 162, 163]. It is also possible to show using triangle inequalities for norms that the lowest possible $l1$-norm for LCU decompositions is $\lambda\geq(E_{\max}-E_{\min})/2$, where $E_{\max/\min}$ are highest/lowest eigenvalues of the Hamiltonian in the entire fermionic Fock space or qubit Hilbert space [161]. Thus, in order to obtain $l1$-norm lower than $(E_{\max}-E_{\min})/2$ one needs to modify the Hamiltonian.
@@
Often in many-body problems, one is interested in eigenstates and eigenenergies in a particular symmetry subspace (e.g., with a fixed number of electrons for molecules). For quantum dynamics, if the initial state lies within a particular symmetry subspace, it continues to evolve there since the Hamiltonian preserves symmetries. In both cases, to reduce $l1$-norm of the LCU decomposition, one can create an effective Hamiltonian that acts identically to the original in the particular symmetry subspace of interest, but has a lower spectral range $(E_{\max}-E_{\min})/2$. In 162, this was done employing the following form for the effective Hamiltonian, $\hat{H}_{\rm eff}=\hat{H}-\hat{O}_{1e}(\hat{N}_{e}-N_{e}\hat{I})+c(\hat{N}_{e}^{2}-N_{e}^{2}\hat{I})$, where $\hat{O}_{1e}$ is an arbitrary one-electron operator, $\hat{N}_{e}$ is the operator of the number of electrons, $N_{e}$ is the number of electrons in the symmetric subspace of interest, and $c$ is a constant. Since finding the true spectral range of $\hat{H}_{\rm eff}$ is difficult, to optimize $\hat{O}_{1e}$ and $c$ in $\hat{H}_{\rm eff}$ one can use the simplest LCU decomposition in terms of Pauli products. $\hat{O}_{1e}$ and $c$ can be optimized to minimize the $l1$-norm of Pauli product decomposition of $\hat{H}_{\rm eff}$.[162] As a direction for further improvement, simulating low-energy states or dynamics by LCU would benefit from either removing from consideration or shifting energies of highly excited states with the right number of electrons so that these states would not limit the $\ell_{1}$-norm of LCU.
@@
Generally, the Pauli products ($\hat{P}_{k}$) are not the only unitaries that one can use for the LCU decomposition. A simple example of improving Pauli products is by grouping them into sets of anti-commuting Pauli products. If $\hat{A}=\sum_{k}c_{k}\hat{P}_{k}$, where $c_{k}$ are real constants and $\{\hat{P}_{k},\hat{P}_{k^{\prime}}\}=0$ then $\hat{A}$ is proportional to the hermitian unitary $\hat{R}=(1/\bar{c})\sum_{k}c_{k}\hat{P}_{k}$, with $\bar{c}=\sqrt{\sum_{k}c_{k}^{2}}$. It is shown in Ref. 161 that such grouping of Pauli products in a qubit Hamiltonian will always reduce the $l1$-norm with respect to the original Pauli product decomposition. There are other forms of unitaries that can be obtained from various fermionic tensor decompositions of the electronic Hamiltonians, for example, tensor hypercontraction or double factorization.[160, 161] To compare different LCU decompositions, one needs to account for three factors: 1) classical cost, 2) $l1$-norm, 3) quantum resources for implementing (e.g., the number of ancilla qubits, T-gate count) of each time step. Note that $l1$-norm mainly affects the size of the time step that one can take, while the third aspect concerns the quantum resources required for each time step.
II.3 Quantum measurement
@@
Many quantum algorithms involve obtaining numerical values of observables, such as energies and dipole moments, as expectation values of associated operators. Valuable use cases could lead to observables on 50 to 100 qubits, which when decomposed into Pauli operators can easily involve millions of terms — for instance, electronic structure Hamiltonians are characterized by approximately $N^{4}$ such Pauli products. Measuring all these Pauli products separately will take a prohibitively long time even on superconducting architectures, which have relatively rapid measurements. Grouping Pauli products is a natural approach to reduce the number of individual operators to measure. Here, we consider various grouping approaches to measure an expectation value of a single operator, like a Hamiltonian, for an arbitrary wavefunction. The generalization of this problem is measuring the expectation values of a set of operators for a set of wavefunctions. Such generalization poses additional optimization issues that can be important for some applications [164], but for the sake of brevity, we focus on the single operator, single wavefunction case for the remainder of this subsection.
@@
All the measurement schemes we describe can be seen as yielding estimators of the Hamiltonian expectation value by combining results of projective measurements of some Pauli operators. One can think of these as either measurements of generic Pauli operators or as measurements of the Ising Hamiltonian terms (i.e., diagonal Pauli operators) with respect to various states related by local basis transformations. Commonly examined physical Hamiltonians, like those for electronic structures, are not readily convertible to Ising form — for instance, through local unitary transformations. Therefore, they must be decomposed into simpler parts. A simple illustration of this idea is an observable $\hat{H}=\sum_{n}\hat{H}_{n}$ that is a sum of Hamiltonian fragments $\hat{H}_{n}$ that can be transformed to the Ising form with some unitary rotations: $\hat{H}_{n}=\hat{U}_{n}^{\dagger}\hat{Z}_{n}\hat{U}_{n}$, where $\hat{Z}_{n}$ is a polynomial of Pauli $\hat{z}$ operators. Then
@@
Within this general scheme, a variety of expectation value estimation methods have been proposed. The methods can be categorized based on several criteria:
@@
    Functional relations: Relations between estimators for measurable operators and that for the total Hamiltonian $\mathit{\mathbb{E}}(\hat{H})=f[\mathit{\mathbb{E}}(\hat{H}_{n})]$ (e.g., linear combinations or more complex functional relations[165]);
@@
One of the estimation techniques that has become popular recently is classical shadow tomography [166]. In the framework described above, the unitary transformations $\hat{U}_{n}$ are chosen randomly from the Clifford group (later, other groups were considered as well [167, 168]). Applying random Clifford transformations ($\{\hat{U}_{k}\}$) as an extension to a state preparation circuit allows one to effectively measure parts of the Hamiltonian that would be transformed to Ising forms $\hat{Z}_{k}$ if $\{\hat{U}_{k}\}$ were applied to the Hamiltonian:
@@
where $\hat{R}_{k}$ is the non-Ising part). This shows that the information about the expectation value of $\hat{H}$ is only obtained from the Ising parts because the expectation value of $\hat{R}_{k}$ on any product state resulting from the wavefunction collapse is zero. This insight spurred various efforts to enhance the sampling convergence by biasing selection of $\{\hat{U}_{k}\}$ to those that increase the norm of $\{\hat{Z}_{k}\}$ parts [169, 170, 171, 172, 173, 174, 175, 176, 177]. A recent paper has benchmarked these randomized measurement methods against a collection of molecular electronic structure Hamiltonians [178].
@@
A more general view of this approach involves the use of adaptive, informationally complete (IC) POVMs [179, 180]. This framework allows the optimization of the measurement for both the observable (e.g. $H$) and the state iteratively while collecting data. Moreover, the same mathematical framework can describe various physical implementations of the measurement [181] and naturally allows for the modelization of noise in the measurement [182]. Finally, it is worth mentioning that classical shadows and IC-POVMs allow for the estimation of multiple observables and for further post-processing, which can include noise mitigation methods [183].
@@
To analyze the efficiency of various estimators one needs to take into account several factors: 1) Classical optimization cost for finding optimal measurable operators; 2) Quantum resource overhead due to the need for extra unitary transformations (e.g., $\hat{U}_{n}$); 3) Measurement overhead due to the number of measurements needed to achieve a desired accuracy in the expectation value. In the near-term, the overhead related to additional unitaries is crucial for overall feasibility. For example, techniques using fermionic $\hat{U}_{n}$’s [184, 185, 186, 187] could be less efficient than those using qubit-based Clifford transformations [188, 189] even though the former require lower numbers of measurements for a given accuracy. Additional options for mitigating circuit and measurement errors become possible using symmetries of measurable fragments. Selecting more symmetric fragments allows one to improve statistics by removing the results that violate expected symmetry constraints (e.g., the number of electrons), which provides an error mitigation protocol [190, 191, 192].
@@
A crucial question for ongoing research is which types of physical Hamiltonians admit feasible measurement schemes in the near term. Spin models typically partition into a few natural measurement bases: for example, a Heisenberg model typically contains terms of the form $XX$, $YY$, $ZZ$, and $Z$, which form three local measurement bases (all $X$, all $Y$, and all $Z$). On the other extreme, molecular electronic structure Hamiltonians have several terms that scale as $\sim N^{4}$ for $N$ orbitals, and it is unclear whether any known measurement scheme can reduce the resulting extremely high measurement counts to something tractable on near-term devices for instances large enough to offer the potential for quantum advantage. Hence an important question is whether there exist intermediate problems where measurement techniques beyond simple grouping as in spin models can be leveraged to achieve tractable measurement counts. For fermionic Hamiltonians, this is closely connected to choices of fermion-to-qubit mapping, as discussed in Subsection II.1. A potential example is the Fermi-Hubbard Hamiltonian, where it is possible to partition the Hamiltonian into 5 operators consisting of commuting Pauli products, and each operator is converted into Ising form by at most one layer of Bell measurement circuit (CNOT + H).[193] By estimating symmetries such as parity, these Bell measurement circuits allow one to mitigate circuit errors.[194]. Another excellent study on the topic as well as other elements of the materials simulation stack is [195].
II.4 Problem-level preprocessing and optimization
@@
Our first example is active space selection (see also Sec. V.3.2). A fundamental problem in chemistry and materials science is the eigenvalue problem, where the eigenstates of the Hamiltonian describing the system are to be solved. Diagonalizing the Hamiltonian of the full system is, however, too expensive and oftentimes unnecessary. In many systems of practical interest, e.g., point defects in wide bandgap semiconductors [196] and catalysts on surfaces or interfaces [197], electronic excited states of extended molecules [198], a certain subset of electrons and orbitals in the molecule or solid is more relevant to the problem than the rest. Therefore, a so-called active space can be defined, leading to an effective Hamiltonian. Solving this effective Hamiltonian (on the quantum computer) is much easier than solving the full Hamiltonian, while the essential properties of the system are retained. Generally speaking, defining active spaces requires identifying a chemically active site on the system. Efforts have been put into developing automated active space selection (see next paragraph), but a universal approach is so far lacking. In practice, prior knowledge of the system studied and the chemical intuition of the researcher is usually relied upon. For example, the spin defect systems, e.g., the nitrogen-vacancy center in diamond and an electronic defect in MgO [199][200, 201, 202] , are heterogeneous materials promising for realizing quantum communication. They have a defect center hosted in a periodic solid, whose local environment is vital to its electronic, optical, and mechanical properties [203]. Therefore, a proper active space should include localized orbitals around the defect center [204, 205]. Other examples are pseudotetrahedral organometallic complexes containing chromium(IV) and aryl ligands which have been experimentally identified as promising molecular qubit candidates [206, 207].
@@
An ab-initio quantum embedding technique, density matrix embedding theory (DMET), [213, 214, 215, 216]and its periodic implementations provide a framework that allows users to effectively reduce an extended system to a finite embedding subspace [217, 218]. The user selects a fragment space characterized by orbitals obtained by using orbital localization schemes. Electrons in such orbitals are generally entangled to the environment, so it is not possible to define a wave function in the fragment orbitals alone. However, the entangled part of the environment can be obtained approximately using the Schmidt decomposition of a single-determinantal whole-system approximate wave function (usually the Hartree–Fock wave function) [215]. This involves the diagonalization of the environmental block ($D_{\text{env}}$) of the one-body reduced density matrix (1-RDM):
@@
Here, $\lambda$ represents a diagonal matrix containing the eigenvalues $\lambda_{i}$ (with $i=0,1,\ldots,N_{\text{env}}$, where $N_{\text{env}}$ indicates the count of environmental orbitals). The columns of the unitary matrix $\mathbf{U}$, corresponding to non-zero and non-two $\lambda_{i}$ values, identify the entangled bath orbitals. Localized fragment orbitals and entangled bath orbitals together define a subspace of the molecule, no larger than twice the size of the original fragment, which is unentangled at the level of the underlying single-determinantal level of theory and within which any quantum chemistry method can be applied by treating all other orbitals as part of a frozen core.
In the case of spin-defects, the fragment active space is usually selected by including orbitals centered around the defect, whereas in the case of gas adsorption, the orbitals centered around the gas molecules and the surface atoms they are bound to are considered. Methods like CAS-DMET and NEVPT2-DMET [200, 201, 202] utilize this framework and provide users the capability of modeling multireference ground and excited states arising from such local perturbations in extended solids with high accuracy. In CAS-DMET, the embedding subspace constructed using equation 7 is subjected to a CASSCF calculation which further reduces the effective active space and accounts for the most relevant static correlation. The remaining dynamic correlation is captured by multireference perturbation theories like NEVPT2[219, 220, 221] as is done in NEVPT2-DMET. A more cost-effective method compared to NEVPT2-DMET to include electron correlation is DME-PDFT [222]. In this approach, the 1- and 2-reduced density matrices (RDMs) generated from a CAS-DMET calculation are used to compute the densities and on-top pair densities, which are then used in a subsequent multiconfiguration pair-density functional theory (MC-PDFT) calculation [223, 224, 225]. MC-PDFT offers a method for integrating multiconfiguration wave function theory and density functional theory, allowing for the comprehensive treatment of both near-degeneracy correlation and dynamic correlation in strongly correlated systems. It proves to be a more cost-effective alternative compared to multireference perturbation theory, multireference configuration interaction, or multireference coupled cluster theory. Moreover, MC-PDFT demonstrates superior accuracy for various properties compared to Kohn–Sham density functional theory. MC-PDFT computes the total energy as a functional of the electron density ($\rho$) and on-top pair density ($\Pi$) obtained from a multiconfiguration wave function. The MC-PDFT energy is expressed as:
@@
where $V_{NN}$ is the nuclear-nuclear repulsion energy, $h_{pq}$ and $g_{pqrs}$ are one- and two-electron integrals, $D_{pq}$ are the elements of the one-electron reduced density matrices (1-RDMs), and the on-top energy $E_{ot}$ is a functional of the density $\rho$ and the on-top pair-density $\Pi$.
It has been used in combination with various multireference wave functions like the generalized active space wave function [226], DMRG [227, 228] and it has been used to compute the total energy of strongly correlated systems starting from the 2-electron reduced density matrix (2-RDM) evaluated on the quantum device.
In the context of DMET, PDFT is especially attractive because the on-top energy is expressed in terms of electronic density on a quadrature grid using semi-local (and therefore linear-scaling) functionals which are agnostic to the origin of the underlying density matrices. This formalism thus models electron correlation of all electrons regardless of any underlying embedding methodology, and DME-PDFT is therefore formally less sensitive to the size of the embedded fragment than methods such as NEVPT2-DMET, which can only model electron correlation within the embedded subspace of fragment and bath orbitals [222].
A promising step towards heterogeneous catalysis applications is the recent study of adsorption with memory-efficient DMET algorithms which, when combined with multireference electronic structure solvers can be extended to the bond-breaking phenomenon on surfaces [229]. A computationally affordable alternative to CASSCF for modeling larger active spaces is the localized active space self-consistent field method (LASSCF) [230, 231]. LASSCF divides a given active space into localized non-interacting subspaces, connected through a mean field, and has the ability to capture strong electron correlation within these subspaces without encountering the computational cost associated with CASSCF.
The LASSCF wavefunction is represented as an antisymmetrized product of the FCI wavefunctions of these individual localized subspaces and the single determinantal wavefunction of the inactive space, which remains delocalized over the entire molecule. This is mathematically represented as:
@@
Although LASSCF treats strong correlations within these localized smaller active spaces, it only accounts for interaction between localized fragments at a mean-field level. Methods such as State Interaction (LASSI) on a classical computer [232] or the variational unitary coupled cluster singles and doubles (UCCSD) on a quantum computer (LAS-UCC) [233] model correlation between these fragments using LASSCF as a reference wave function. In the LAS-UCC method, the multiconfigurational subspace wavefunction $\Psi_{A_{K}}$ for each fragment is prepared using either Quantum Phase Estimation (QPE) or Direct Initialization (DI), as outlined in D’cunha et al. 2023 [234]. The inter-fragment correlations are then captured through a standard Variational Quantum Eigensolver (VQE) optimization process using the Unitary Coupled Cluster (UCC) ansatz.

### v2

@@ III.2.2 Linear Combination of Unitaries
Another approach to the simulation of $e^{-iHt}$ involves block-encoding the Hamiltonian as a part of a larger unitary, $U(H)$. One of the most typical block-encodings is accomplished by writing $H$ as a linear combination of unitaries, $H=\sum_{n}c_{n}U_{n}$, and building $U(H)$ using $U_{P}=\sum_{n}\mathcal{P}_{n}\otimes U_{n}$, where $\{\mathcal{P}_{n}\}$ is a complete set of orthogonal projectors, $\mathcal{P}_{n}\mathcal{P}_{n^{\prime}}=\mathcal{P}_{n}\delta_{nn^{\prime}}$ built in the ancilla qubit space. The computational cost of the algorithms based on linear combination of unitaries (LCU) depends on the $\ell_{1}$-norm of $c_{n}$ coefficient vector, $\lambda=\sum_{n}|c_{n}|$ [160]. Thus, much effort has been directed to reducing the $\ell_{1}$-norm by various selections of $U_{n}$ operators [161, 162, 163, 164]. It is also possible to show using triangle inequalities for norms that the lowest possible $\ell_{1}$-norm for LCU decompositions is $\lambda\geq(E_{\max}-E_{\min})/2$, where $E_{\max/\min}$ are highest/lowest eigenvalues of the Hamiltonian in the entire fermionic Fock space or qubit Hilbert space [162]. Thus, in order to obtain $\ell_{1}$-norm lower than $(E_{\max}-E_{\min})/2$ one needs to modify the Hamiltonian.
@@
Often in many-body problems, one is interested in eigenstates and eigenenergies in a particular symmetry subspace (e.g., with a fixed number of electrons for molecules). For quantum dynamics, if the initial state lies within a particular symmetry subspace, it continues to evolve there since the Hamiltonian preserves symmetries. In both cases, to reduce $\ell_{1}$-norm of the LCU decomposition, one can create an effective Hamiltonian that acts identically to the original in the particular symmetry subspace of interest, but has a lower spectral range, $(E_{\max}-E_{\min})/2$. In [163], this was done by employing the form of the effective Hamiltonian, $\hat{H}_{\rm eff}=\hat{H}-\hat{O}_{1e}(\hat{N}_{e}-N_{e}\hat{I})+c(\hat{N}_{e}^{2}-N_{e}^{2}\hat{I})$, where $\hat{O}_{1e}$ is an arbitrary one-electron operator, $\hat{N}_{e}$ is the operator of the number of electrons, $N_{e}$ is the number of electrons in the symmetric subspace of interest, and $c$ is a constant. Since finding the true spectral range of $\hat{H}_{\rm eff}$ is difficult, to optimize $\hat{O}_{1e}$ and $c$ in $\hat{H}_{\rm eff}$ one can use the simplest LCU decomposition in terms of Pauli products. $\hat{O}_{1e}$ and $c$ can be optimized to minimize the $\ell_{1}$-norm of Pauli product decomposition of $\hat{H}_{\rm eff}$. As a direction for further improvement, simulating low-energy states or dynamics via LCU would benefit from either removing from consideration or shifting energies of highly excited states with the right number of electrons so that these states would not limit the $\ell_{1}$-norm of LCU.
@@
Generally, the Pauli products ($\hat{P}_{k}$) are not the only unitaries that one can use for the LCU decomposition. A simple example of improving Pauli products is by grouping them into sets of anti-commuting Pauli products. If $\hat{A}=\sum_{k}c_{k}\hat{P}_{k}$, where $c_{k}$ are real constants and $\{\hat{P}_{k},\hat{P}_{k^{\prime}}\}=0$ then $\hat{A}$ is proportional to the Hermitian unitary $\hat{R}=(1/\bar{c})\sum_{k}c_{k}\hat{P}_{k}$, with $\bar{c}=\sqrt{\sum_{k}c_{k}^{2}}$. It is shown in Ref. [162] that such grouping of Pauli products in a qubit Hamiltonian will always reduce the $\ell_{1}$-norm with respect to the original Pauli product decomposition. There are other forms of unitaries that can be obtained from various fermionic tensor decompositions of the electronic Hamiltonians, for example, tensor hypercontraction or double factorization [161, 162]. To compare different LCU decompositions, one needs to account for three factors: 1) classical cost, 2) $\ell_{1}$-norm, 3) quantum resources for implementing (e.g., the number of ancilla qubits, T-gate count) of each time step. Note that $\ell_{1}$-norm mainly affects the size of the time step that one can take, while the third aspect concerns the quantum resources required for each time step.
III.3 Quantum measurement
@@
Many quantum algorithms involve obtaining numerical values of observables, such as energies and dipole moments, as expectation values of their associated operators. Valuable use cases could lead to observables on 50 to 100 qubits, which when decomposed into Pauli operators can easily involve millions of terms; for instance, electronic structure Hamiltonians are characterized by approximately $N^{4}$ such Pauli products. Measuring all these Pauli products separately will take a prohibitively long time even on superconducting architectures, which have relatively rapid measurements. Grouping Pauli products is a natural approach to reduce the number of individual operators to measure. Here, we consider various grouping approaches to measure an expectation value of a single operator, like a Hamiltonian, for an arbitrary wavefunction. The generalization of this problem is measuring the expectation values of a set of operators for a set of wavefunctions. Such generalization poses additional optimization issues that can be important for some applications [165], but for the sake of brevity, we focus on the single operator, single wavefunction case for the remainder of this subsection.
@@
All measurement schemes we describe can be seen as yielding estimators of the Hamiltonian expectation value by combining results of projective measurements of some Pauli operators. One can think of these as either measurements of generic Pauli operators or as measurements of the Ising Hamiltonian terms (i.e., diagonal Pauli operators) with respect to various states related by local basis transformations. Commonly examined physical Hamiltonians, like those for electronic structures, are not readily convertible to Ising form; for instance, through local unitary transformations. Therefore, they must be decomposed into simpler parts. A simple illustration of this idea is an observable $\hat{H}=\sum_{n}\hat{H}_{n}$ that is a sum of Hamiltonian fragments $\hat{H}_{n}$ that can be transformed to the Ising form with some unitary rotations: $\hat{H}_{n}=\hat{U}_{n}^{\dagger}\hat{Z}_{n}\hat{U}_{n}$, where $\hat{Z}_{n}$ is a polynomial of Pauli $\hat{z}$ operators. Then
@@
Within this general scheme, a variety of expectation value estimation methods have been proposed. The methods can be categorized based on the following criteria:
@@
    Functional relations: Relations between estimators for measurable operators and that for the total Hamiltonian, $\mathit{\mathbb{E}}(\hat{H})=f[\mathit{\mathbb{E}}(\hat{H}_{n})]$ (e.g., linear combinations or more complex functional relations[166]);
@@
One of the estimation techniques that has become popular recently is classical shadow tomography [167]. In the framework described above, the unitary transformations $\hat{U}_{n}$ are chosen randomly from the Clifford group (later, other groups were considered as well [168, 169]). Applying random Clifford transformations, $\{\hat{U}_{k}\}$, as an extension to a state preparation circuit allows one to effectively measure parts of the Hamiltonian that would be transformed to Ising forms $\hat{Z}_{k}$ if $\{\hat{U}_{k}\}$ were applied to the Hamiltonian:
@@
where $\hat{R}_{k}$ is the non-Ising part). This shows that the information about the expectation value of $\hat{H}$ is only obtained from the Ising parts because the expectation value of $\hat{R}_{k}$ on any product state resulting from the wavefunction collapse is zero. This insight spurred various efforts to enhance the sampling convergence by biasing selection of $\{\hat{U}_{k}\}$ to those that increase the norm of $\{\hat{Z}_{k}\}$ parts [170, 171, 172, 173, 174, 175, 176, 177, 178]. A recent paper [179] has benchmarked these randomized measurement methods against a collection of molecular electronic structure Hamiltonians.
@@
A more general view of this approach involves the use of adaptive, informationally incomplete (IC) Positive Operator Valued Measure (POVM) [180, 181]. This framework allows the optimization of the measurement for both the observable (e.g., $H$) and the state iteratively while collecting data. Moreover, the same mathematical framework can describe various physical implementations of the measurement [182] and naturally allows for the modelization of noise in the measurement [183]. Finally, it is worth mentioning that classical shadows and IC-POVMs allow for the estimation of multiple observables and for further post-processing, which can include noise mitigation methods [184].
@@
To analyze the efficiency of various estimators one needs to take into account several factors: 1) classical optimization cost for finding optimal measurable operators; 2) quantum resource overhead due to the need for extra unitary transformations (e.g., $\hat{U}_{n}$); and 3) measurement overhead due to the number of measurements needed to achieve a desired accuracy in the expectation value. In the near-term, the overhead related to additional unitaries is crucial for overall feasibility. For example, techniques using fermionic $\hat{U}_{n}$ [185, 186, 187, 188] could be less efficient than those using qubit-based Clifford transformations [189, 190] even though the former require lower numbers of measurements for a given accuracy. Additional options for mitigating circuit and measurement errors become possible using symmetries of measurable fragments. Selecting more symmetric fragments allows one to improve statistics by removing the results that violate expected symmetry constraints – e.g., the number of electrons – which provides an error mitigation protocol [191, 192, 193].
@@
A crucial question for ongoing research is which types of physical Hamiltonians admit feasible measurement schemes in the near term. Spin models typically partition into a few natural measurement bases; e.g., a Heisenberg model typically contains terms of the form $XX$, $YY$, $ZZ$, and $Z$, which form three local measurement bases (all $X$, all $Y$, and all $Z$). On the other extreme, molecular electronic structure Hamiltonians have several terms that scale $\sim N^{4}$ for $N$ orbitals. It is unclear whether any known measurement scheme can reduce the resulting extremely large measurement counts to something tractable on near-term devices for instances large enough to offer the potential for quantum advantage. Hence an important question is whether there exist intermediate problems where measurement techniques beyond simple grouping – as in spin models – can be leveraged to achieve tractable measurement counts. For fermionic Hamiltonians, this is closely connected to choices of fermion-to-qubit mapping, as discussed in Sec. III.1. A potential example is the Fermi-Hubbard Hamiltonian, where it is possible to partition the Hamiltonian into 5 operators consisting of commuting Pauli products, and each operator is converted into Ising form by at most one layer of Bell measurement circuit (Controlled NOT gate (CNOT) + H) [194]. By estimating symmetries such as parity, these Bell measurement circuits allow one to mitigate circuit errors [195]. Another excellent study on the topic as well as other elements of the materials simulation stack can be found in Ref. [196].
III.4 Problem-level preprocessing and optimization
@@
Our first example is active space selection (see also Sec. VII.3.2). A fundamental problem in chemistry and materials science is the eigenvalue problem, where the eigenstates of the Hamiltonian describing the system are to be solved. Diagonalizing the Hamiltonian of the full system is, however, too expensive and oftentimes unnecessary. In many systems of practical interest, e.g., point defects in wide bandgap semiconductors [197] and catalysts on surfaces or interfaces [198], electronic excited states of extended molecules [199], a certain subset of electrons and orbitals in the molecule or solid is more relevant to the problem than the rest. Therefore, a so-called active space can be defined, leading to an effective Hamiltonian. Solving this effective Hamiltonian (on a quantum computer) is much easier than solving the full Hamiltonian, while the essential properties of the system are retained. Generally speaking, defining active spaces requires identifying a chemically active site on the system. Efforts have been put into developing automated active space selection (see next paragraph), but a universal approach is so far lacking. In practice, prior knowledge of the system studied and the chemical intuition of the researcher is usually relied upon. For example, spin defect systems, e.g., the nitrogen-vacancy center in diamond and an electronic defect in MgO [200, 201, 202, 203], are heterogeneous materials promising for the realization of quantum communication. They have a defect center hosted in a periodic solid, whose local environment is vital to its electronic, optical, and mechanical properties [204]. Therefore, a proper active space should include localized orbitals around the defect center [205, 206]. Other examples are pseudotetrahedral organometallic complexes containing chromium(IV) and aryl ligands which have been experimentally identified as promising molecular qubit candidates [207, 208].
@@
An ab-initio quantum embedding technique, density matrix embedding theory (DMET)  [214, 215, 216, 217], and its periodic implementations provide a framework that allows users to effectively reduce an extended system to a finite embedding subspace [218, 219]. The user selects a fragment space characterized by orbitals obtained by using orbital localization schemes. Electrons in such orbitals are generally entangled to the environment, so it is not possible to define a wave function in the fragment orbitals alone. However, the entangled part of the environment can be obtained approximately using the Schmidt decomposition of a single-determinant whole-system approximate wave function (usually the Hartree–Fock wave function) [216]. This involves the diagonalization of the environmental block, $D_{\text{env}}$, of the one-body reduced density matrix (RDM):
@@
Here, $\lambda$ represents a diagonal matrix containing the eigenvalues $\lambda_{i}$ for $i=0,1,\ldots,N_{\text{env}}$, and where $N_{\text{env}}$ indicates the count of environmental orbitals. The columns of the unitary matrix $\mathbf{U}$, corresponding to non-zero and non-two $\lambda_{i}$ values, identify the entangled bath orbitals. Localized fragment orbitals and entangled bath orbitals together define a subspace of the molecule, no larger than twice the size of the original fragment, which is unentangled at the level of the underlying single-determinantal level of theory and within which any quantum chemistry method can be applied by treating all other orbitals as part of a frozen core.
In the case of spin-defects, the fragment active space is usually selected by including orbitals centered around the defect, whereas in the case of gas adsorption, the orbitals centered around the gas molecules and the surface atoms they are bound to are considered. Methods like CAS-DMET and NEVPT2-DMET [201, 202, 203] utilize this framework and provide users the capability of modeling multireference ground and excited states arising from such local perturbations in extended solids with high accuracy. In CAS-DMET, the embedding subspace constructed using Eqn. (7) is subjected to a CASSCF calculation which further reduces the effective active space and accounts for the most relevant static correlation. The remaining dynamic correlation is captured by multireference perturbation theories like NEVPT2 [220, 221, 222] as is done in NEVPT2-DMET.
A more cost-effective method compared to NEVPT2-DMET to include electron correlation is DME-PDFT [223]. In this approach, the 1- and 2-RDMs generated from a CAS-DMET calculation are used to compute the densities and on-top pair densities, which are then used in a subsequent multiconfiguration pair-density functional theory (MC-PDFT) calculation [224, 225, 226]. MC-PDFT offers a method for integrating multiconfiguration wave function theory and density functional theory, allowing for the comprehensive treatment of both near-degeneracy correlation and dynamic correlation in strongly correlated systems. It proves to be a more cost-effective alternative to multireference perturbation theory, multireference configuration interaction, or multireference coupled cluster theory. Moreover, MC-PDFT demonstrates superior accuracy for various properties compared to Kohn–Sham density functional theory. MC-PDFT computes the total energy as a functional of the electron density ($\rho$) and on-top pair density ($\Pi$) obtained from a multiconfiguration wave function.
The MC-PDFT energy is expressed as:
@@
where $V_{NN}$ is the nuclear-nuclear repulsion energy, $h_{pq}$ and $g_{pqrs}$ are one- and two-electron integrals, $D_{pq}$ are the elements of the one-electron RDM, and the on-top energy $E_{ot}$ is a functional of the density $\rho$ and the on-top pair-density $\Pi$. It has been used in combination with various multireference wave functions like the generalized active space wave function [227], DMRG [228, 229] and it has been used to compute the total energy of strongly correlated systems starting from the 2-electron RDM evaluated on the quantum device.
In the context of DMET, PDFT is especially attractive because the on-top energy is expressed in terms of electronic density on a quadrature grid using semi-local (and therefore linear-scaling) functionals which are agnostic to the origin of the underlying density matrices. This formalism thus models electron correlation of all electrons regardless of any underlying embedding methodology. Therefore, DME-PDFT is formally less sensitive to the size of the embedded fragment than methods such as NEVPT2-DMET, which can only model electron correlation within the embedded subspace of fragment and bath orbitals [223].
A promising step towards heterogeneous catalysis applications is the recent study of adsorption with memory-efficient DMET algorithms which, when combined with multireference electronic structure solvers, can be extended to the bond-breaking phenomenon on surfaces [230]. A computationally affordable alternative to CASSCF for modeling larger active spaces is the localized active space self-consistent field (LASSCF) method [231, 232]. LASSCF divides a given active space into localized non-interacting subspaces, connected through a mean field, and has the ability to capture strong electron correlation within these subspaces without encountering the computational cost associated with CASSCF.
The LASSCF wavefunction is represented as an antisymmetrized product of the Full Configuration Interaction (FCI) wavefunctions of these individual localized subspaces and the single determinantal wavefunction of the inactive space, which remains delocalized over the entire molecule. This is mathematically represented as:
@@
Although LASSCF treats strong correlations within these localized smaller active spaces, it only accounts for interactions between localized fragments at a mean-field level. Methods such as State Interaction (LASSI) on a classical computer [233] or the variational unitary coupled cluster singles and doubles (UCCSD) on a quantum computer (LAS-UCC) [234] model correlation between these fragments using LASSCF as a reference wave function. In the LAS-UCC method, the multiconfigurational subspace wavefunction $\Psi_{A_{K}}$ for each fragment is prepared using either QPE or Direct Initialization (DI), as outlined by D’cunha et al. [235]. The inter-fragment correlations are then captured through a standard VQE optimization process using the Unitary Coupled Cluster (UCC) ansatz.